makaylab:

In 25 hour, Matt can type 45 page. What is Matt's rate in pages per hour? Drag and drop the correct value into the box. Response area pages per hour Answers:3 1/8 1/2 8/25 2

makaylab:

25/is 2/5 and 45 is 4/5

ramen:

First find the rate matt can type in 1 hour.

makaylab:

how tho @ramen

ramen:

Wait you said that it isn't 25 hours it is 2/5 of an hour and not 45 pages but 4/5 of a page?

makaylab:

yes

makaylab:

@ramen

ramen:

ok

makaylab:

kk

ramen:

Well 1/5 of an hour = 12 minutes because 60/5=12 you understand?

makaylab:

no

ramen:

ok so because 2/5 is a fraction, that means that it is 2 out of the 5 equal parts in 1 hour. 60 minutes = 1 hour, so 60/5=12 because there are 5 equal parts of 12 minutes in an hour.

makaylab:

bu did you read the answers @ramen

ramen:

I'm only done with part 1 of the work

makaylab:

ok

ramen:

So I hope you understood what I just explained?

ramen:

Now that we got the basics of fractions down

makaylab:

i got it ramen

ramen:

You would do (4/5)/(2/5) which is = to 4/5*5/2 because the fraction switches when it is multiplied.

ramen:

What do you get when you do this?

ramen:

That is your answer.

makaylab:

2

ramen:

Yes good job

ramen:

You can medal me if you think I helped.